OpenStudy (anonymous):

How to Solve : 64329 is divided by a certain number. While dividing the numbers 175,114, and 213 appear as three successive remainders. The divisor is (1) 184 (2) 224 (3) 234 (4) 296

ganeshie8 (ganeshie8):

say the required number is \(n\)

ganeshie8 (ganeshie8):

then, \(n\) must satisfy : \(643 = p n + 175 \) \(1752 = qn + 114\) \(1149 = rn + 213\)

ganeshie8 (ganeshie8):

if that makes any sense...

OpenStudy (anonymous):

ofcourse that do.... thanks bro...

OpenStudy (anonymous):

468, 1638, 936... i got , how does that relate to question. Wondering wer u get 1752 from?

ganeshie8 (ganeshie8):

how did u get 1638 ?

OpenStudy (anonymous):

1752-114=1638

ganeshie8 (ganeshie8):

oh you used those equations ok :)

ganeshie8 (ganeshie8):

to see from 1752 is coming from, u need to think of "long division"

ganeshie8 (ganeshie8):